# BoxingGym: Benchmarking Progress in Automated Experimental Design and Model Discovery

### v1

@@ 4 Experiments with Baseline Agents
LLM Agent. We prompt an LLM, GPT-4o, to interact with our environment directly, purely through natural language, without any additional functionalities (see Fig. 2, see App. for details).
Box’s Apprentice. Box’s Apprentice is an agent that augments the statistical modelling capabilities of the language model (e.g., GPT-4o) by allowing it to write and fit a generative model to observed data and use them to make predictions [21] (see App. C for implementation details). This allows Box’s Apprentice to build explicit representations of the underlying phenomena in each environment. In particular, for the model discovery experiments, we ask the LM agent to write a pymc program given the collected datapoints after 10 experiments. We then fit this generative model and provide the implementation of the model and its inferred parameters to the scientist producing the explanation for the novice agent. For the experimental design experiments, we ask the LM agent to write a pymc program given the collected datapoints so far. The agent then uses this model to guide its subsequent experiments.
Experiment Setup. For each environment, we run the agents for 5 independent trials. At each step, the agent chooses to perform an experiment, by specifying a design, and observes the outcome. After a fixed number of steps (0, 1, 3, 5, 7, 10), we evaluate the agent’s performance using the metrics described earlier § 3.2. The values reported in Tab. 1 and Tab. 2 are averaged across the 5 runs and over 10 evaluation points. We also explore a prior vs no prior conditions. In the prior condition, we give the LM full context about the problem domain (e.g., you are observing how participants balance delayed vs immediate rewards) . In the no prior condition, we remove this context and describe the setting in a domain-agnostic way (e.g., you receive a tuple of three values).
@@ 4.1 Experimental Design Evaluation
Figure 3: Standardized errors compared. We plot the standardized errors for the two agents, gpt-4o (solid line) and Box’s Apprentice (dashed line) across three domains: Peregrines (left), Hyperbolic Discounting (center) and IRT (right). Error bars are 95% CIs across 5 runs.
@@
To evaluate the agents’ performance, we first assess their ability to gather valuable information through their experiment selection and then measure how effectively they use this information to predict the environment. The Expected Information Regret (EI Regret), a key metric, compares the Expected Information Gain (EIG) (§ 3.2.1) of the agent’s chosen experiments to the maximum EIG achievable from 100 random experiments. Lower EI Regret indicates more informative experiment selection. To verify if agents can use observational data to make predictions, we measure their standardized prediction error about the environment before and after experimentation.
We present our findings [^3] We omit the predatory-prey and Emotions domains for Box’s Apprentice, since GPT-4o could not reliably produce pymc programs in Tab. 1, Tab. 2 and Fig. 3 (see Tab. 3 for EI regrets). We observe that in a few environments, the LLM agent possesses accurate prior knowledge, and in these cases, experimentation can occasionally lead to a decline in performance as the language model “over-corrects” based on a limited number of observations (see hyperbolic discounting; Tab. 1). For instance, in the hyperbolic discounting environment (Fig. 3, center), the error after experimentation is higher than before conducting experiments. The LM agent excels at integrating observations for the Peregrines and Predator-Prey environments, as indicated by the negative standardized error. However, for most other environments, its performance does not significantly improve with experimentation and remains suboptimal, as evidenced by the positive standardized error.
Although Box’s Apprentice can propose and fit a pymc model to the observed data, it fails to systematically improve its predictions consistently after additional observations. For example, it outperforms the pure LLM agent in the death process environment (-1.02 compared to 0.46 for the LLM agent), but is worse at Peregrines (0.04 compared to -0.65 for the LLM agent). We observe that one common failure mode is that, because the number of datapoints is relatively small, Box’s Apprentice has a consistent bias towards simple functional forms. For example, to model the population count of Peregrines, Box’s Apprentice starts with a simple linear model, and after collecting more datapoints, opts to parameterize the Poisson rate as a linear function of the year, which is a restrictive model. In some cases, Box’s Apprentice identifies a program that is very close to the ground truth program; for example, in the IRT domain, Box’s Apprentice proposes a logistic model that incorporates student abilities and question difficulties as features. This is also the case for the Location Finding domain; Box’s Apprentice comes up with inverse square functional form (App. Fig. 8). However, as its high prediction error indicates, it is not able to strategically collect datapoints that are useful for fitting an accurate probabilistic model in the low data regime. This is consistent with our EIG results in App. Tab. 3 showing that the scientist agent’s experiments attain high regret relative to a random baseline.
4.2 Model Discovery Evaluation via Communication
@@
Next, we evaluate the agents’ ability to build and communicate models that capture the underlying phenomena in each environment. To test this, we have the agents interact with the environment for 10 steps (scientist phase) and then generate a natural language explanation of their findings. We then provide this explanation to a novice agent, which must make predictions about the environment without any direct interaction (novice phase by using the explanation from the scientist; § 3.2.2). We consider two conditions: 1) Both the scientist and novice agents receive a brief prior description of the environment. 2) Neither the scientist nor the novice receives any prior information. The scientist’s prediction after 10 observations (Error After Experiments) acts as a weak positive control, and the scientist’s predictions without any observations (Error Before Experiments) acts as a negative control. Ideally, if the scientist’s explanation is effective, the novice’s error should be lower than the negative control and approach the positive control. For Box’s Apprentice, we provide the scientist with the statistical model and ask it to use both the statistical model and its experiments to produce an explanation.
The results in Tab. 1 and Tab. 2 show that the explanations generated by both agents generally improve the novice’s predictions, but both agents struggle across different environments. Despite having access to an explicit generative model, Box’s apprentice does not always produce a better explanation. In some cases, this is because a generative model does not enable the scientist to draw any additional useful conclusions as the agent’s prior knowledge for the domain is already substantial. For example, in the IRT domain, the agent summarizes the parameters of the explicit statistical model as revealing that question difficulty and student ability are both related to a student’s likelihood of answering a question correctly (see App. Fig. 10). We find that a purely natural language based agent already understands this about the environment and can also produce a similar explanation, as indicated by the language agent’s performance on IRT compared to Box’s Apprentice (App. Fig. 9). Surprisingly, for IRT, we find that having to communicate an explanation to the novice agent forces the LLM agent to generalize its findings, allowing the novice agent to achieve a significantly lower error compared to the scientist (-0.28 compared to 0.00 for the LLM agent). In general, there is a gap between the novice’s performance and the positive control, showing that learning from an agent-generated explanation is not as informative as performing interactive experiments directly with the environment. As a surprising exception, for IRT, we find that having to communicate an explanation to the novice agent forces the LLM agent to generalize its findings, allowing the novice agent to achieve a significantly lower error compared to the scientist (-0.28 compared to 0.00).
@@ 5 Discussion
We introduced BoxingGym , a benchmark for measuring experimental design coupled with model discovery with a language interface. BoxingGym includes 10 environments, each based on real-world generative models, that allow us to systematically compare different language-based agents. To evaluate experimental design capabilities, we used information gain metrics to approximate the effectiveness of an agent’s active experimentation. For model discovery and communication, we introduced a novel metric based on the agent’s ability to explain its discovered model to a novice agent, who then makes predictions about the environment. Our experiments with two types of agents—a standard language model and a language model augmented with statistical modeling—demonstrate this framework’s potential to drive progress. Results with our baseline agents reveal significant challenges: proposing hypotheses from few observations, designing experiments from hypotheses, and incorporating new observations with prior knowledge to revise them.
While BoxingGym is a valuable testbed, it has limitations. The environments in BoxingGym offer pre-defined experimental paradigms, whereas real-world science involves designing the experimental method itself; deciding variables to measure, conditions to compare, and controlling for confounds. Future research could develop environments requiring agents to design experiments from scratch [12]. Additionally, BoxingGym does not consider the time and cost of experiments. However, scientists work within resource constraints and decide the number of experiments based on time, cost, and information gain. Future versions could incorporate these costs, allowing agents to choose the number of experiments or evaluate informativeness normalized by cost (e.g., [47]). Lastly, while BoxingGym covers various scientific domains, it is still limited compared to the full breadth of scientific inquiry. Expanding the benchmark to include more diverse fields, such as biology, chemistry, and social science, would provide a more comprehensive test.
One unique aspect of BoxingGym is the inclusion of environments that involve human behavior, such as the Moral Machines and Emotions environments. These domains require the agent to reason about psychological factors and social dynamics, which are crucial for many real-world scientific applications. Future work could incorporate more sophisticated simulations of human participants [3, 1, 46, 35, 36], to create more realistic simulations of experiments in psychology and the social sciences.
Our experiments showcased the potential of interfaces, such as access to statistical modelling libraries, that augment language models’ reasoning capabilities for experimental design and discovery. However, there is still much room for improvement in the design of interfaces that help a model understand an environment and the data produced better. Future research in designing agents could explore new interfaces that leverage data visualization techniques, strategic simulations based on prior theories, model validation [22], and even web-based research strategies, to enhance the model’s ability to help guide experimentation and discovery [21, 51].

### v2

@@ 4 Experiments
LLM Agent.
We consider 6 LLMs, GPT-4o [39], Claude-3.7-sonnet [3], Qwen-2.5-32b-instruct, Qwen-2.5-7b-instruct [55], and reasoning variants OpenThinker-32b, and OpenThinker-7b [51]; the reasoning variants are finetuned on math and coding task. We prompt these models to interact with our environment, purely through natural language, without additional tools (see Fig. 2, see App. B for details).
Box’s Apprentice.
The apprentice agent augments language models by enabling them to implement generative models of observed data. For model discovery, the agent writes a pymc program [27] after 10 experiments, which is then fit and provided to the scientist explaining findings to the novice. For experimental design, the agent creates and uses these models to guide subsequent experiments.
Experiment Setup. For each environment, we run the agents for 5 independent trials. At each step, the agent chooses to perform an experiment, by specifying a design, and observes the outcome. After a fixed number of steps (0, 1, 3, 5, 7, 10), we evaluate the agent’s performance using the metrics described earlier § 3.2. The performance of models is averaged across 5 runs and over 10 evaluation points. We also explore a prior vs no prior condition to investigate whether domain knowledge helps or hinders scientific discovery. In the prior condition, we give the LM full context about the problem domain (e.g., “you are observing how participants balance delayed vs immediate rewards”), simulating scientists with background knowledge. In the no prior condition, we remove this context and describe the setting in a domain-agnostic way (e.g., “you receive a tuple of three values”), resembling reasoning from raw observations without preconceptions. This tests whether prior knowledge scaffolds discovery or creates biases that constrain exploration.
@@ 4.1 Experimental Design Evaluation
Figure 3: Normalized Error Compared across Models. (a) Comparison of the normalized errors for different LLMs with or without prior information included in the prompt. (b) Comparison of reasoning models (OpenThinker) and instruct models (Qwen) across environments. Error bars are the standard error across 5 runs.
Figure 4: Normalized Errors Over Number of Observations. Normalized errors for the LLM agent with gpt-4o with prior information (solid blue) and without prior information (dotted yellow) across three domains: Population Growth Dynamics (left), IRT (center) and Hyperbolic Discounting (right). Error bars are the standard error across 5 runs.
Setup.
@@
To evaluate the agents’ performance, we first assess their ability to gather valuable information through their experiment selection and then measure how effectively they use this information to predict the environment. The Expected Information Regret (EI Regret) compares the Expected Information Gain (EIG) (§ 3.2.1) of the agent’s chosen experiments to the maximum EIG achievable from 100 random experiments. Lower EI Regret indicates more informative experiment selection.
Prior information does not improve performance.
We find that models often perform better when given no prior information after 10 experiments (Fig. 3a). In some cases, this is because the LLM makes an overly strong assumption about the environment (e.g., the signal decay is symmetric around the origin) and does not revise the assumption after more experiments; this is consistent with findings reported by Li et al. [27]. In other cases, such as the hyperbolic discounting environment (Fig. 4, right), the model overfits to limited observations.
More experiments generally lead to better predictions.
We plot the learning trajectories for three environments in (Fig. 4). The agent’s average prediction error decreases as it performs more experiments. The Hyperbolic Temporal Discounting environments shows an unexpected trends where more experiments actually increases error. This may again be related to how prior knowledge interferes with effective learning from data.
Models Improve with Scale.
Larger models consistently outperform their smaller counterparts within the same model family. Both OpenThinker-32B and Qwen2.5-32B demonstrate significantly better performance than their respective 7B variants across environments (Fig. 3a), highlighting the benefits of scale for experimental design tasks.
Instruction-Tuned Models outperform Reasoning Models.
Surprisingly, the instruction-tuned Qwen2.5 models outperform the reasoning-focused OpenThinker models (Fig. 3b). This may be because OpenThinker models are finetuned to perform well on a relatively narrow set of verifiable problems in math and code, while instruction-tuned models retain broader capabilities that could be useful for experimental design.
Models performance varies substantially across environments.
Models show varying performance across different environments (Fig. 3b). Performance is strongest on environments like population growth dynamics and death process, where the LM agent achieves negative standardized error, indicating that the LM successfully leveraged information gained through experimentation. However, in environments like hyperbolic discounting, performance is low even after experimentation, suggesting that some domains are inherently more challenging for current models.
EIG Regret reveals relationship between experimental design and prediction.
Our EIG regret analysis (Fig. 5b) provides insight into the relationship between two key components of scientific reasoning: designing informative experiments and making accurate predictions from collected data. GPT-4o achieves both the lowest EIG regret and strong predictive performance across several environments, suggesting these capabilities can be aligned. However, the varying performance of other models is informative — for instance, Qwen-32B shows higher EIG regret despite good predictive performance in some domains, indicating that while these abilities may be related, excellence in prediction doesn’t automatically translate to optimal experimental design.
LLMs cannot always optimally leverage statistical models.
While Box’s Apprentice can propose and fit explicit statistical models to observed data, it does not consistently improve over the non-augmented LLM (GPT-4o) (Fig. 5a) From qualitative analysis of the models, we find that Box’s Apprentice tends to favor overly simple functional forms due to limited data, such as using linear approximations for inherently nonlinear phenomena.
Figure 5: (a) Comparison of the Box’s Apprentice with an LLM agent. (b) EIG Regret scores for six large language models, with lower values indicating better performance.
4.2 Evaluating Model Discovery via Communication
Figure 6: Evaluation of Model Discovery via Communication. (a) Comparison of the standardized error of the Novice (gpt-4o) with different Scientist models. (b) Comparison of errors made by the Novice and the Scientist (both models are gpt-4o). Error bars are standard error.
Setup.
@@ 4.2 Evaluating Model Discovery via Communication
Next, we evaluate the agents’ ability to build and communicate models that capture the underlying phenomena in each environment. To test this, we have the agents interact with the environment for 10 steps (scientist phase) and then generate a natural language explanation of their findings. We then provide this explanation to a novice agent, which must make predictions about the environment without any direct interaction (novice phase by using the explanation from the scientist; § 3.2.2). The novice agent is always gpt-4o. The scientist’s prediction after 10 observations (Error After Experiments) acts as a weak positive control. Ideally, if the scientist’s explanation is effective, the novice’s error should approach the positive control.
Explanations improve with scale.
Larger models generally produce more effective explanations, as evidenced by better novice performance when using explanations from 32B variants compared to 7B models (Fig. 6a). This suggests that increased model scale improves not just experimentation but also the ability to distill and communicate findings.
Explanations are not as good as experiments.
As expected, novice agents perform worse than scientists who directly interacted with the environment (Fig. 6b). The gap suggests that current explanation methods do not fully capture the knowledge gained through experimentation.
Explanations are more helpful for some environments.
However, the effectiveness of explanations varies substantially across domains (Fig. 6b). For instance, explanations are helpful for animal growth, but struggle with complex domains like moral judgments. This variation likely reflects the complexity of different domains and the current limitations of language models in capturing and communicating certain types of patterns.
@@ 5 Discussion
We introduced BoxingGym , a benchmark measuring language-based agents’ capabilities in experimental design and model discovery across 10 real-world-based environments. We evaluated experimental design using information gain metrics and developed a novel model discovery metric based on an agent’s ability to explain its model to a novice agent. Our evaluation across multiple model scales (7B-32B parameters) shows that while larger and closed-source models generally perform better, fundamental challenges persist. Neither domain-specific prior knowledge nor statistical modeling capabilities consistently improved performance. Some environments yielded strong results with larger models, while others remained challenging for all approaches. BoxingGym has limitations: it uses pre-defined experimental paradigms rather than requiring design from scratch [14], ignores resource constraints, and covers limited scientific domains. Future work should address these limitations by incorporating experiment design from scratch, resource constraints, and more diverse fields [23]. We could also expand the human behavior environments (Moral Machines, Emotions) with more sophisticated participant simulations [4, 1, 50, 40, 41]. While our experiments demonstrated potential for interfaces that augment language models’ scientific reasoning capabilities, future research should explore data visualization, model validation [28], and web-based research strategies to enhance experimental guidance and discovery.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We describe the design of our benchmark accurately, summarize results with different models.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See discussion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No proofs or new theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, further, all our code, results and scripts are available on github.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All the code is accessible on the github.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe this in detail in experimental setup and have the full specification in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report statistical significance in all our results…
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See appendix section B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Single blind submission and we follow the code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: We don’t discuss these as there are no direct negative societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Not relevant for the paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All models have been cited appropriately. The papers that inspired the environments have been credited too.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We add documentation to the BoxingGym code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human participants were recruited.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Paper does not use human participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: None of the core methods used LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.